British Exporter 1 million Danish Krona
England Denmark
Interest 2% 8%
Borrow 3% 9%
Spot 8.300-01 K/L
Forward 9.000-01 K/L
Hedge with Forwards
Hedge with Money Market Hedge
Which do you prefer?
Answer #1

Part A:

Amount thriugh Fwd Rate = 1000000 / 9.001

= GBP 111,098.77

Part B:

Amount through Money Market Hedge:

Borrow in Danish krona whose future value equals to 1M in a year.

Amount borrowed = 1000000 / 1.09

= 917,431.19

Convert into GBP using spot rate

= 917431.19 / 8.301

= GBP 110,520.56

Deposit the amount for 1 Year in London and realize the maturity

= GBP 110,520.56 * 1.02

= 112,730.97

Money Market Hedge is suggested as more GBP return is there.

Pls comment, if any further assistance is required.